# HiSAC: High-Resolution Sensing with Multiband Communication Signals

### v1

@@ Abstract.
Integrated Sensing And Communication (ISAC) systems are expected to perform accurate radar sensing while having minimal impact on communication. Ideally, sensing should only reuse communication resources, especially for spectrum which is contended by many applications. However, this poses a great challenge in that communication systems often operate on narrow subbands with low sensing resolution. Combining contiguous subbands has shown significant resolution gain in active localization. However, multiband ISAC remains unexplored due to communication subbands being highly sparse (non-contiguous) and affected by phase offsets that prevent their aggregation (incoherent). To tackle these problems, we design HiSAC, the first multiband ISAC system that combines diverse subbands across a wide frequency range to achieve super-resolved passive ranging. To solve the non-contiguity and incoherence of subbands, HiSAC combines them progressively, exploiting an anchor propagation path between transmitter and receiver in an optimization problem to achieve phase coherence. HiSAC fully reuses pilot signals in communication systems, it applies to different frequencies and can combine diverse technologies, e.g., 5G-NR and WiGig. We implement HiSAC on an experimental platform in the millimeter-wave unlicensed band and test it on objects and humans. Our results show it enhances the sensing resolution by up to 20 times compared to single-band processing while occupying the same spectrum.
@@ 1. Introduction
Endowing wireless communication systems with radar sensing capabilities is one of the key objectives of 3GPP Sixth Generation (6G) and future Wi-Fi (Lu et al., 2024). In recent years, so-called ISAC systems have enabled a wide range of applications from multitarget tracking (Wu et al., 2020; Pegoraro et al., 2023), person identification (Pegoraro et al., 2022; Yang et al., 2020), activity recognition (Meneghello et al., 2022; Lin et al., 2023), vital signs monitoring (Zeng et al., 2020), pose estimation (Ji et al., 2024; Jiang et al., 2020), and object imaging (Zhang et al., 2021b; Yao et al., 2024).
@@
Motivation. A fundamental trade-off in ISAC systems is to achieve high sensing resolution, i.e., the capability of distinguishing multiple closely located targets, and accuracy with minimal impact on the primary communication functionality. Ideally, sensing should be performed by fully reusing resources available to the communication system in time and frequency. Particular attention has to be put on spectrum, which is becoming more and more scarce due to the ubiquitous applications of Radio Frequency (RF) transmissions (Testolina et al., 2024). Indeed, ISAC benefits from using a large bandwidth since this is inversely proportional to the ranging resolution, i.e., the minimum signal propagation distance under which two targets can not be distinguished. However, the bandwidth available to existing communication systems is insufficient to achieve the desired cm-level ranging resolution in 6G. Even wideband Fifth Generation-New Radio (5G-NR) channels and IEEE 802.11ay (WiGig) in the Millimeter-Wave (mmWave) band can at most reach 37 cm and 17 cm resolution with 400 MHz and 1.76 GHz bandwidth, respectively. Such resolution can be improved by applying super-resolution algorithms based on MUltiple SIgnal Classification (MUSIC) or Compressed Sensing (CS), e.g., (Samczyński et al., 2021; Kotaru et al., 2015; Khalilsarai et al., 2020), but the bandwidth limitation remains.
@@
A possible solution to enhance the resolution is to combine multiple communication frequency bands to increase the sensing bandwidth. This approach has been attempted for active localization in Orthogonal Frequency Division Multiplexing (OFDM) systems (where the user carries a communication device) (Kotaru et al., 2015; Xiong et al., 2015; Wan et al., 2023; Noschese et al., 2020), radar (Cuomo et al., 1999; Xiong et al., 2017; Zhang et al., 2017), and recently for sub-6 GHz Wi-Fi sensing in (Wang et al., 2024). However, several limitations make the above methods unsuitable for ISAC. On the one hand, active localization approaches exploit either contiguous or closely-spaced subbands, which may not be available in ISAC since the spectrum is contended by a plethora of services and contains frequency gaps. Moreover, they can count on a collaborative localized device, which simplifies the problem since synchronization errors that prevent accurate delay measurements can be resolved via handshaking (Khalilsarai et al., 2020). On the other hand, radar methods use dedicated waveforms, optimized for sensing purposes, and relatively wide subbands to be combined, which significantly simplifies the problem. Lastly, (Wang et al., 2024) employs a neural network model to overcome the above limitations in Wi-Fi, but this ties the system to the specific frequency band, communication technology (Wi-Fi), and hardware used to collect the training data. Conversely, we aim to develop a system that seamlessly adapts to OFDM, Single Carrier (SC), cellular, or Wi-Fi.
@@
Challenges. Designing such a system presents several open challenges. First, one must tackle the non-contiguity of communication systems subbands, which may include gaps of several hundreds of MHz or even GHz. Second, the different subbands are affected by time-varying and unknown timing, frequency, and phase offsets that prevent the coherent combination of the Channel Frequency Response (CFR) estimated by the communication protocol over different ISAC receivers (Zhang et al., 2022; Wu et al., 2024). Although the compensation of timing and frequency offsets in ISAC systems has been widely studied (Zhang et al., 2020; Pegoraro et al., 2024; Zhu et al., 2018; Ni et al., 2021; Li et al., 2022), phase synchronization is not well investigated since it is not needed in typical sensing tasks such as target tracking and Doppler estimation. On the contrary, achieving phase coherence is a strict requirement to combine multiple subbands over a wide frequency range. Third, communication subbands are relatively narrow with respect to the total bandwidth required to achieve high resolution. This makes it difficult to model individual subbands as they contain an insufficient number of frequency samples. Conversely, reconstructing the CFR over the total bandwidth entails huge computational complexity due to the high number of subcarriers. Finally, the designed method should generalize to different communication systems (OFDM vs. SC), protocols (5G-NR vs. Wi-Fi), and channel representations (e.g., CFR vs. Channel Impulse Response (CIR)).
@@
We implement HiSAC on a Radio Frequency System on a Chip (RFSoC) platform in the mmWave unlicensed band (58-64 GHz). We demonstrate that HiSAC achieves a few-cm ranging resolution on metal and human targets, giving a 3 to 20 times improvement over baseline methods. Moreover, it works in mono-/bi-static configurations on typical multiband systems employing Carrier Aggregation (CA), Bandwidth Part (BWP), and it is robust to target motion.
@@
3. Our approach entails zero additional overhead on communication and seamlessly integrates with multiband communication systems that adopt CA or BWP.
@@ 2. Preliminaries
In this section, we provide useful background on the applicability of HiSAC in ISAC systems, ranging resolution, and phase incoherence due to phase offsets.
@@ 2.1. HiSAC use cases in ISAC
Carrier aggregation. CA is a function implemented in the Radio Access Network (RAN) and User Equipments of 5G-NR mobile wireless networks. CA combines multiple frequency allocations (carriers) at different radio cells to boost the data rate of the connection (Li et al., 2023). A set of serving cells is identified that contributes to the aggregation, which can take place within the same frequency band (intra-band), e.g., within the sub-$6$ GHz range, or across multiple bands (inter-band), e.g., across Frequency Range (FR2222 and FR $3$. Exploiting the channel estimation process carried out on each frequency band for sensing would greatly enhance the available bandwidth and, as a result, the ranging resolution.
Bandwidth part. BWP is a mechanism to split the radio channel available to a cell into multiple segments (parts), which can then be used to allocate different regions of the spectrum to different applications (or UEs(Lin et al., 2021). Only one part can be active at a time, and SSBs can be transmitted in each part to synchronize to different UEs. Given that SSBs have a relatively narrow bandwidth, it is appealing to develop a system that can combine the SSBs transmitted by one or more radio cells to perform accurate mono-static ranging by exploiting the total frequency aperture over a wider bandwidth. Note that, BWP poses the additional requirement that the system must be able to operate with CFR estimates that are not collected simultaneously in all the subbands.
@@ 2.2. Delay and ranging resolution
Improving ranging resolution is a challenging task since the bi-static angle depends on the location of the TX, RX, and target, and increasing the bandwidth is not viable in ISAC systems since it is pre-determined by the communication protocol. Super-resolution approaches have been proposed that exploit subspace-based methods, e.g., MUSIC (Kotaru et al., 2015), or CS algorithms (Khalilsarai et al., 2019), exploiting assumptions on the structure of the channel (e.g., sparsity). However, none of these methods can drastically improve the ranging resolution, and limitations due to the narrow bandwidth remain.
@@ 2.3. Phase-incoherence among subbands
In a wireless communication system, the clock signal of each node is generated from a Local Oscillator (LO). The different LOs are asynchronous, meaning that, due to hardware non-idealities, they are subject to time-varying relative drifts from their nominal oscillating frequencies (Zhang et al., 2021a). In addition, their initial phase is random. This introduces unwanted offsets in the received signals that are specific to each TX-RX pair. These can be categorized into Timing Offset (TO), Carrier Frequency Offset (CFO), and Random Phase Offset (RPO(Wu et al., 2024).
@@
When multiple ISAC systems in different frequency bands are considered, a direct combination of their CFRs is infeasible due to the presence of the above offsets. Indeed, this causes the phases of the CFRs in the different bands to be misaligned at the different RXs, preventing the construction of a common model spanning the full frequency band. Several approaches have been proposed to tackle TO and CFO in ISAC systems (Zhang et al., 2020; Pegoraro et al., 2024; Zhu et al., 2018; Ni et al., 2021; Zeng et al., 2019, 2020; Li et al., 2022; Chen et al., 2023; Meneghello et al., 2022). However, none of these tackles phase synchronization by also eliminating the RPO, which is essential for multiband CFR combination.
@@ 3.1. Non-coherent subsystems and subbands
Consider a wide frequency band with bandwidth $B$ and starting frequency $f_{0}$, which we denote as the full band of interest, as shown in Fig. 2. We call $\Delta_{f}$ the spacing of the frequency samples of the considered CFR (subcarriers). $\Delta_{f}$ corresponds to the subcarrier spacing in OFDM systems or to the Discrete Fourier Transform (DFT) samples spacing used in SC systems. The total number of virtual subcarriers in the full band is $K=B/\Delta_{f}$, indexed by $k=0,\dots,K-1$. We use the term virtual to highlight that not all the subcarriers are used for communication, which is carried out on a subset of the full band spectrum. In the following, we assume the same subcarrier spacing $\Delta_{f}$ is shared by all the considered subbands, which can be achieved using interpolation or downsampling.
@@ 3.2. Multiband channel model
In this section, we present the multiband CFR model. We consider a time-varying multipath channel with $L$ propagation paths, where $t$ is used to denote time. We denote by $\tau_{l}(t)$ and $\alpha_{l}(t)$ the delay of the $l$-th channel path due to propagation and its complex amplitude at time $t$, accounting for the combined effect of the propagation loss and the target’s scattering phase (Richards et al., 2010).

### v2

@@ Abstract.
Integrated Sensing And Communication (ISAC) systems are expected to perform accurate radar sensing while having minimal impact on communication. Ideally, sensing should only reuse communication resources, especially for spectrum which is contended by many applications. However, this poses a great challenge in that communication systems often operate on narrow subbands with low sensing resolution. Combining contiguous subbands has shown significant resolution gain in active localization. However, multiband ISAC remains unexplored due to communication subbands being highly sparse (non-contiguous) and affected by phase offsets that prevent their aggregation (incoherent). To tackle these problems, we design HiSAC, the first multiband ISAC system that combines diverse subbands across a wide frequency range to achieve super-resolved passive ranging. To solve the non-contiguity and incoherence of subbands, HiSAC combines them progressively, exploiting an anchor propagation path between transmitter and receiver in an optimization problem to achieve phase coherence. HiSAC fully reuses pilot signals in communication systems, applies to different frequencies, and can combine diverse technologies, e.g., 5G-NR and WiGig. We implement HiSAC on an experimental platform in the millimeter-wave unlicensed band and test it on objects and humans. Our results show it enhances the sensing resolution by up to 20 times compared to single-band processing while occupying the same spectrum.
Integrated sensing and communications, human sensing, multiband, super-resolution, Wi-Fi sensing, 5G.
journalyear: 2024copyright: rightsretainedconference: The 22nd ACM Conference on Embedded Networked Sensor Systems; November 4–7, 2024; Hangzhou, Chinabooktitle: The 22nd ACM Conference on Embedded Networked Sensor Systems (SENSYS ’24), November 4–7, 2024, Hangzhou, Chinadoi: 10.1145/3666025.3699357isbn: 979-8-4007-0697-4/24/11ccs: Hardware Signal processing systemsccs: Human-centered computing Ubiquitous and mobile computingccs: Applied computing Telecommunications
@@ 1. Introduction
Endowing wireless communication systems with radar sensing capabilities is one of the key objectives of 3GPP Sixth Generation (6G) and future Wi-Fi (Lu et al., 2024). In recent years, so-called ISAC systems have enabled a wide range of applications from multitarget tracking (Wu et al., 2020; Pegoraro et al., 2024b), person identification (Pegoraro et al., 2022; Yang et al., 2020), activity recognition (Meneghello et al., 2023; Lin et al., 2023), vital signs monitoring (Zeng et al., 2020), pose estimation (Ji et al., 2024; Jiang et al., 2020), and object imaging (Zhang et al., 2021c; Yao et al., 2024).
@@
Motivation. A fundamental trade-off in ISAC systems is to achieve high sensing resolution, i.e., the capability of distinguishing multiple closely located targets, and accuracy with minimal impact on the primary communication functionality. Ideally, sensing should be performed by fully reusing resources available to the communication system in time and frequency. Particular attention has to be put on spectrum, which is becoming increasingly scarce due to the ubiquitous applications of Radio Frequency (RF) transmissions (Testolina et al., 2024). Indeed, ISAC benefits from using a large bandwidth since this is inversely proportional to the ranging resolution, i.e., the minimum signal propagation distance below which two targets can not be distinguished. Note that enhancing ranging resolution is also highly beneficial to advanced sensing applications such as human respiration monitoring. Although small chest displacements can be monitored even with a narrowband signal (Zhang et al., 2021a), using the carrier phase, separating the respiration of multiple closely-located subjects with low ranging resolution is extremely challenging and prone to errors
However, the bandwidth available to existing communication systems is insufficient to achieve the desired cm-level ranging resolution in 6G. Even wideband Fifth Generation-New Radio (5G-NR) channels and IEEE 802.11ay (WiGig) in the Millimeter Wave (mmWave) band can at most reach 37 cm and 17 cm resolution with 400 MHz and 1.76 GHz bandwidth, respectively. Such resolution can be improved by applying super-resolution algorithms based on MUltiple SIgnal Classification (MUSIC) or compressed sensing, e.g., (Samczyński et al., 2021; Kotaru et al., 2015; Khalilsarai et al., 2020), but the bandwidth limitation remains.
@@
A possible solution to enhance the resolution is to combine multiple communication frequency bands to increase the sensing bandwidth. This approach has been attempted for active localization in Orthogonal Frequency Division Multiplexing (OFDM) systems (where the user carries a communication device) (Kotaru et al., 2015; Xiong et al., 2015; Wan et al., 2023; Noschese et al., 2020), radar (Cuomo et al., 1999; Xiong et al., 2017; Zhang et al., 2017), and recently for sub-6 GHz Wi-Fi sensing in (Li et al., 2024). However, several limitations make the above methods unsuitable for ISAC. On the one hand, active localization approaches exploit either contiguous or closely-spaced subbands, which may not be available in ISAC since the spectrum is contended by a plethora of services and contains frequency gaps. Moreover, they can count on a collaborative localized device, which simplifies the problem since synchronization errors that prevent accurate delay measurements can be resolved via handshaking (Khalilsarai et al., 2020). On the other hand, radar methods use dedicated waveforms, optimized for sensing purposes, and relatively wide subbands to be combined, which significantly simplifies the problem. Lastly, (Li et al., 2024) employs a neural network model to overcome the above limitations in Wi-Fi, but this ties the system to the specific frequency band, communication technology (Wi-Fi), and hardware used to collect the training data. Conversely, we aim to develop a system that seamlessly adapts to different modulation types such as 5G-NR OFDM and IEEE 802.11ay Single Carrier (SC).
@@
Challenges. Designing such a system presents several open challenges. First, one must tackle the non-contiguity of communication systems subbands, which may include gaps of several hundreds of MHz or even GHz. Second, the different subbands are affected by time-varying and unknown timing, frequency, and phase offsets that prevent the coherent combination of the Channel Frequency Response (CFR) estimated by the communication protocol over different ISAC receivers (Zhang et al., 2022; Wu et al., 2024). Although the compensation of timing and frequency offsets in ISAC systems has been widely studied (Zhang et al., 2020; Pegoraro et al., 2024a; Zhu et al., 2018; Ni et al., 2021; Li et al., 2022), phase synchronization is not well investigated since it is not needed in typical sensing tasks such as target tracking and Doppler estimation. On the contrary, achieving phase coherence is a strict requirement to combine multiple subbands over a wide frequency range. Third, communication subbands are relatively narrow with respect to the total bandwidth required to achieve high resolution. This makes it difficult to model individual subbands as they contain insufficient frequency samples. Conversely, reconstructing the CFR over the total bandwidth entails huge computational complexity due to the high number of subcarriers. Finally, the designed method should generalize to different communication systems (OFDM vs. SC), protocols (5G-NR vs. Wi-Fi), and channel representations (e.g., CFR vs. Channel Impulse Response (CIR)).
@@
We implement HiSAC on a Radio Frequency System on a Chip (RFSoC) platform in the mmWave unlicensed band (58-64 GHz). We demonstrate that HiSAC achieves a few-cm ranging resolution on metal and human targets, giving a 3 to 20 times improvement over baseline methods. Moreover, it works in mono-/bi-static configurations on typical multiband systems employing carrier aggregation, bandwidth part, and it is robust to target motion.
@@
3. Our approach entails zero additional overhead on communication and seamlessly integrates with multiband communication systems that adopt carrier aggregation or bandwidth part.
@@ 2. Preliminaries and motivation
In this section, we provide useful background on the applicability of HiSAC in ISAC systems, ranging resolution, and phase incoherence due to phase offsets. Then, we highlight the novelty of our approach with respect to existing active multiband systems for localization.
@@ 2.1. HiSAC use cases in ISAC
Carrier aggregation. Carrier aggregation is a function implemented in the Radio Access Network (RAN) and User Equipments of 5G-NR mobile wireless networks. carrier aggregation combines multiple frequency allocations (carriers) at different radio cells to boost the data rate of the connection (Li et al., 2023). A set of serving cells is identified that contributes to the aggregation, which can take place within the same frequency band (intra-band), e.g., within the sub-$6$ GHz range, or across multiple bands (inter-band), e.g., across Frequency Range (FR2222 and FR $3$. Exploiting the channel estimation process carried out on each frequency band for sensing would greatly enhance the available bandwidth and, as a result, the ranging resolution.
Bandwidth part. Bandwidth part is a mechanism to split the radio channel available to a cell into multiple segments (parts), which can then be used to allocate different regions of the spectrum to different applications (or UEs(Lin et al., 2021). Only one part can be active at a time, and SSBs can be transmitted in each part to synchronize to different UEs. Given that SSBs have a relatively narrow bandwidth, it is appealing to develop a system that can combine the SSBs transmitted by one or more radio cells to perform accurate mono-static ranging by exploiting the total frequency aperture over a wider bandwidth. Note that, bandwidth part poses the additional requirement that the system must be able to operate with CFR estimates that are not collected simultaneously in all the subbands.
@@ 2.2. Delay and ranging resolution
Improving ranging resolution is a challenging task since the bi-static angle depends on the location of the TX, RX, and target, and increasing the bandwidth is not viable in ISAC systems since it is pre-determined by the communication protocol. Super-resolution approaches have been proposed that exploit subspace-based methods, e.g., MUSIC (Kotaru et al., 2015), or compressed sensing algorithms (Khalilsarai et al., 2019), exploiting assumptions on the structure of the channel (e.g., sparsity). However, none of these methods can drastically improve the ranging resolution, and limitations due to the narrow bandwidth remain.
@@ 2.3. Phase-incoherence among subbands
In a wireless communication system, the clock signal of each node is generated from a Local Oscillator (LO). The different LOs are asynchronous, meaning that, due to hardware non-idealities, they are subject to time-varying relative drifts from their nominal oscillating frequencies (Zhang et al., 2021b). In addition, their initial phase is random. This introduces unwanted offsets in the received signals that are specific to each TX-RX pair. These can be categorized into Timing Offset (TO), Carrier Frequency Offset (CFO), and Random Phase Offset (RPO(Wu et al., 2024).
@@
When multiple ISAC systems in different frequency bands are considered, a direct combination of their CFRs is infeasible due to the presence of the above offsets. Indeed, this causes the phases of the CFRs in the different bands to be misaligned at the different RXs, preventing the construction of a common model spanning the full frequency band. Several approaches have been proposed to tackle TO and CFO in ISAC systems (Zhang et al., 2020; Pegoraro et al., 2024a; Zhu et al., 2018; Ni et al., 2021; Zeng et al., 2019, 2020; Li et al., 2022; Chen et al., 2023; Meneghello et al., 2023). However, none of these tackles phase synchronization by also eliminating the RPO, which is essential for multiband CFR combination.
2.4. Innovation over multiband localization
Exploiting multiple frequency bands for localizing an active receiver device has been widely studied. HiSAC instead tackles the different problem of passive ranging, in which the target of the localization is an object or person, not necessarily equipped with a radio device.
Two main aspects prevent the application of existing methods for active localization to the passive ISAC scenario.
First, active localization targets the time-of-arrival estimation of the Line-of-Sight (LOS) path between TX and RX, as done by Chronos (Vasisht et al., 2016) and SpotFi (Kotaru et al., 2015), among others (Bansal et al., 2021; Khalilsarai et al., 2020). Hence, multipath reflections are regarded as a nuisance and have to be compensated for. Conversely, our work focuses on obtaining the propagation delays of reflections on passive objects that are causing the multipath effect. These are significantly weaker than the LOS, due to the longer propagation distance and scattering on the target. The applicability of active localization methods for paths other than the LOS has not been demonstrated, making them unsuitable for radar-like sensing, which is the objective of HiSAC.
Second, existing multiband localization methods combine overlapping, contiguous, or relatively close frequency subbands (Kotaru et al., 2015; Xie et al., 2015; Xiong et al., 2015). This is due to the limitations of existing algorithms to achieve phase coherence, which do not scale well to wide frequency bands, and the high computational complexity of compressed sensing. Due to these limitations, existing methods can not be applied to ISAC, where subbands may be widely separated in frequency as discussed in Section 2.1, with a low ratio of measured frequencies over the total bandwidth aperture. We demonstrate instead that HiSAC can accurately perform ranging with as low as $1/6$ measured frequencies ratio (see Section 6.2). This is enabled by exploiting the channel sparsity, originally applying compressed sensing in two incremental steps to reduce complexity, and to the increased robustness of HiSAC’s phase offsets compensation method. Thanks to these innovations, HiSAC is also the first multiband system that demonstrates cross-technology sensing capabilities, combining OFDM and SC channel estimates over wide bands.
[…]
@@ 3.1. Non-coherent subsystems and subbands
Consider a wide frequency band with bandwidth $B$ and starting frequency $f_{0}$, denoted as the full band of interest, as shown in Fig. 2. The full band represents the total frequency aperture of HiSAC, which gives the ideal delay resolution of the system. In practice, the full band may represent the total aperture of carrier aggregation or bandwidth part systems, spanning from the frequency of the first estimated subcarrier in the channel response to the last one, including frequency gaps.
We call $\Delta_{f}$ the spacing of the frequency samples of the considered CFR (subcarriers). $\Delta_{f}$ corresponds to the subcarrier spacing in OFDM systems or to the Discrete Fourier Transform (DFT) samples spacing used in SC systems. The total number of virtual subcarriers in the full band is $K=B/\Delta_{f}$, indexed by $k=0,\dots,K-1$. We use the term virtual to highlight that not all the subcarriers are used for communication, which is carried out on a subset of the full band spectrum. In the following, we assume that all the considered subbands share the same subcarrier spacing $\Delta_{f}$, which can be achieved using interpolation or downsampling.
@@
Note that HiSAC handles subsystems and subbands with different bandwidths, which is critical to seamlessly integrate it into communication systems operating in diverse frequency bands.
@@ 3.2. Multiband channel model
In this section, we present the multiband CFR model. We consider a time-varying multipath channel with $L$ propagation paths, where $t$ is used to denote time. We denote by $\tau_{l}(t)$ and $\alpha_{l}(t)$ the delay of the $l$-th channel path due to propagation and its complex amplitude at time $t$, accounting for the combined effect of the propagation loss and the target’s scattering phase (Richards et al., 2010). Tab. 1 summarizes the notation used in the system model.
@@
Start frequency $f_{0}$ No. avail. subc. subsys. $i$ $M_{i}$
Full bandwidth $B$ TO subsystem $i$ $\tau_{o,i}$
Subcarrier spacing $\Delta_{f}$ CFO subsystem $i$ $f_{o,i}$
Total no. subcarriers $K$ RPO subsystem $i$ $\varphi_{o,i}$
No. of subsystems $C$ PO subsystem $i$ $\phi_{o,i}$
No. subcarriers subsys. $i$ $K_{i}$ Total delay resolution $\Delta\tau$
No. subc. subsys. $i$ subb. $s$ $K_{i,s}$ Number of paths $L$
Start frequency subsys. $i$ $f_{i}$ Complex gain path $l$ $\alpha_{l}$
Bandwidth subsystem $i$ $B_{i}$ Delay path $l$ $\tau_{l}$
Subbands set subsystem $i$ $\mathcal{S}_{i}$ Start idx. subsys. $i$ subb. $s$ $k_{i,s}$
No. subbands subsystem $i$ $S_{i}$ $k$-th subc. in total CFR $H_{k}$
No. available subcarriers $M$ $k$-th subc. subsys. $i$ subb. $s$ $H_{i,s,k}$
Table 1. Summary of the notation.